# Improved Prediction Of Ligand-Protein Binding Affinities By Meta-Modeling Ho-Joon Lee1,2,*#, Prashant S. Emani3,*# and Mark B. Gerstein3,4,5,6,7#

### v2

@@ Introduction
In this work, we aim to improve ligand-protein binding affinity prediction by combining both classical scoring functions, which are built upon physical force fields, and sequence-based DL models into a single integrative framework of meta-models (also called ensemble, fusion, or hybrid models). Meta-models are a type of ensemble learning by combining diverse algorithms and have been successful in machine or statistical learning for improved results with more robustness and less bias than individual algorithms (known as stacked generalization or super learning) 74, 75. Meta-models, termed blending in associated publications, were also very successful in addressing the Netflix Prize challenges 76-79. In predicting ligand-protein binding affinity, ensemble models of multiple 3D CNNs have been shown to outperform single CNNs and previous models 50. On the other hand, theory-driven physical models and data-driven ML or statistical models are complementary to each other. The former offers physical interpretability and constraints, whereas the latter is better at complex pattern discovery from massive data. For this reason, hybrid approaches have been favored in recent years 72, 73, 80-83. In general, the intuition behind meta-modeling is that different solutions to a problem could be modeling non-overlapping aspects of the ground truth. These different solutions may be trained on data with their own systematic biases, or might have architectures that access some limited subspace of the possible model space. When combined, an ensemble of such solutions might serve to better represent the ground truth. We have designed a modeling framework that allows the inclusion of structure-based classical scoring functions and sequence-based DL models, as well as explicit inclusion of covariates such as physicochemical features of ligand, proteins, or ligand-protein complexes (Fig. 1 and Supplementary Fig. S1). In this study, we first focus on 2 empirical scoring functions (ESFs), SMINA and Vinardo (implemented in smina) 20 together with consensus docking strategies based on experimental or predicted ligand poses. For DL models, we use sequence-based DL models of 10 different architectures with multiple training strategies, employing an open-source DL library for drug-target interaction prediction called DeepPurpose 57 and two ligand-protein binding databases, BindingDB 84 and PDBbind. We investigate both pre-trained models from the DeepPurpose library and our own fine-tuned or de novo-trained models using PDBbind, which is the main data used by the ESF models. In addition, by merging all prediction scores from repeated cross-validations for all DL models under study, we investigate top principal components from principal component analysis (PCA) as additional prediction scores. Finally, by taking all prediction scores from both ESF and DL models as well as molecular weights of ligands, we build meta-models using 4 machine learning algorithms and 11 feature sets as our final tools (44 in total) for ligand-protein binding affinity prediction.
@@ Methods
We calculate in two ways, depending on the choice of docking poses: (1) Experimental pose-RMSD filtering, where the RMSD is calculated for a ligand pose predicted by a docking method relative to the experimental structure of the ligand (when available); and (2) Consensus pose-RMSD filtering, where the RMSD is calculated between the docking poses for the same ligand-target pair as predicted by two different docking scoring methods, Vinardo and SMINA. Details of these filtering approaches are provided in Supporting Information. We also chose a cutoff of 3 Angstroms as representing a reasonably good quality match between poses based on our manual inspection of RMSD distributions (Supplementary Fig. S2A). This is a more lenient cutoff relative to other analyses 20, 23 that suggest a 2-Angstrom cutoff. Our approach is a compromise between the need for higher quality matches, and the need to have sufficient numbers of data points to train our models as stricter cutoffs lead to smaller dataset sizes.
@@
Virtual Screening Benchmark
@@ Results
As our first group of DL models, we investigated either pre-trained (i.e., published) or de novo-trained models using the BindingDB (BDB) or PDBbind-RefinedSet (PDBb) data (model class D1, D2, or D3, respectively) (Table 2). Based on mean-squared-error (MSE), we selected the following top 6 out of the 12 de novo-trained models for subsequent analyses, which show comparable performance to the DeepPurpose pre-trained models with MSE < 0.6: CNN-CNN, Daylight-AAC, Morgan-AAC, Morgan-CNN, MPNN-CNN, and Transformer-CNN (Supplementary Fig. S3). As the overlap between the published and de novo-trained models is only two, CNN-CNN and MPNN-CNN, this expanded the family of BDB-trained base models to a total of 10 with competitive performances in this study (Table 2). In terms of mean cross-validation (from 10x 5-fold CV), the best performing BDB-pre-trained, BDB-de-novo-trained, and PDBb-trained models were Daylight-AAC, Daylight-CNN, and CNN-AAC (Table 2). The model average Pearson correlation coefficients were 0.468, 0.475, and 0.746 for the CoreSet, respectively (Fig. 3A; D1, D2, and D3). The two BDB-trained models perform worse than the docking tools (Fig. 2). Full performance results are provided in Supplementary Tables S2 and S3. We note that the PDBb-trained models performed poorly for BDB-Kd predictions with no correlations (Supplementary Fig. S4).
@@
As our second group of models, we investigated fine-tuned models using PDBb for both BDB-pre-trained and BDB-de-novo-trained models (D1F and D2F, respectively). In terms of mean cross-validation (from 10x 5-fold CV), the best performing PDBb-finetuned models for BDB-pre-trained and BDB-de-novo-trained models were CNN-CNN and MPNN-AAC. The model average Pearson correlation coefficients were 0.716 and 0.749 for the CoreSet, respectively (Fig. 3A). Full performance results are provided in Supplementary Table S3. As for BDB-Kd prediction, all PDBb-finetuned models maintained good performances with Pearson correlations of 0.6–0.7, compared to predictions by the BDB-trained models (Supplementary Figs. S4 and S5).
@@
As our third or final group of DL models, we investigated PCA projections of all predictions from those fine-tuned models for the BDB-pre-trained (D1F) or BDB-de-novo-trained (D2F) in the second model group (6*5*10=300 D1F or D2F model instances from 10x 5-fold CV for each model architecture) or those PDBb-trained models (D3) in the first model group (10*5*10=500 model instances). The first principal component (PC1) shows the best performance in all PCA, Pearson correlation coefficients for D1F, D2F, and D3 being 0.715, 0.749, and 0.747, respectively (Fig. 3A; D1FP, D2FP, and D3P). We also performed PCA on the merged set of all predictions from D1F, D2F, and D3 (300+300+500=1100 model instances), the Pearson correlation coefficient for PC1 being 0.744 (Fig. 3A; DAP). As expected, the PC1 is highly correlated with the mean prediction with the largest Pearson correlation coefficient of > 0.99 in all cases, while the second most correlated PC was not the second PC (PC2). By looking at performances of all cross-validated fine-tuned or PCA models, we observe that PC1 and D3 models tend to perform better than D1 or D2 models for the CoreSet in terms of Pearson correlation coefficients or mean squared errors (Fig. 3B).
@@
Given that we have investigated the 38 DL models (6 D1, 6 D2, 10 D3, 6 D1F, 6 D2F, 1 D1FP, 1 D2FP, 1 D3P, and 1 DAP) as well as the 4 docking models, we first compared their predictions by Pearson correlation values among themselves. Fig. 4A shows a heatmap of Pearson correlation values for all pairs of predictions for the CoreSet. As expected, high correlations (i.e., similar predictions) are observed among (1) the 4 docking models and (2) the fine-tuned or PC1 models (D1F, D2F, D1FP, D2FP, D3P). The BindingDB-based models (D1 and D2) are less correlated among themselves. Correlations between the experimental binding affinity and those predictions show that the top model is D1F|Daylight_AAC (pre-trained and fine-tuned) and the next top 3 models are PC1 models (Fig. 4B). We also performed another comparison analysis by prediction errors or deviations from the experimental binding affinity (Supplementary Table S3). Figs. 4C and 4D show a heatmap and a box-whisker plot of prediction deviations or absolute errors for the 42 models. The fine-tuned or PC1 models tend to show lower errors than the BindingDB models, which in turn tend to show lower errors than the docking models. The top 5 models in terms of median errors are the 4 PC1 models and D2F|MPNN_Transformer (de novo-trained and fine-tuned). Similar but more pronounced correlations and deviations are observed for the training set, as expected (Supplementary Fig. S6).
@@
To further parse the differences between the outputs of the prediction tools, we assessed whether certain properties of the proteins or ligands are over- or under-represented in the complexes which were: (1) Best modeled by empirical docking; (2) Best modeled by deep learning; (3) Best modeled by meta-modeling. We also explored the top 50 complexes as predicted by the XGBoost Meta-model for ED-A-P to see if the performance of the meta-models relative to the experimental data was contingent on protein or ligand properties. We extracted protein features from the UniProt database and ligand features from the SwissADME database. We tested the significance of the occurrence of certain annotations (using Fisher’s exact test) and of the distributions of certain quantitative features (using a two-sided Wilcoxon test) for each of the four subsets of complexes. Specifically, we considered whether several UniProt features were annotated more or less frequently, and whether certain Gene Ontology (GO) 92, domain and pathway annotations occurred with different probabilities in the subsets. We also compared the distributions of the ligand SwissADME features. The results (Supplementary Table S9) indicate that there are practically no significant (nominal p-value ) protein features prioritized in the four cases considered. This may be due to the fact that several proteins (and their associated UniProt IDs) are duplicated in the CoreSet. The number of reviewed proteins with UniProtIDs drops from 243 to 60 after removal of all duplicates. The highly significant results with nominal p-values (both Wilcoxon and t-tests) are the distributions of several ligand features including MW, lipophilicity, and solubility measures in the structures for which the empirical docking tools perform the best. There is also some hint of significance for MW, the “#Rotatable bonds” and “#H-bond acceptors” in the structures for which the meta-models perform the best.
@@
Overall, we conclude that the family of our models are competitive to several state-of-the-art structure-based models and also useful for a virtual screening application.
@@ Discussion
Overall, our framework corroborates existing work on ensemble learning, stacked generalization, super learner, and consensus-based improvements in predictions. Our machine-learning-based meta-models allow for the flexibility of adding or removing any component tools, physicochemical properties, or molecular descriptors as part of feature engineering. Models trained on multiple datasets (e.g. BindingDB and PDBbind) and on multiple data types (e.g. sequence and/or structure) may be unified to reduce individual biases and improve predictions. Furthermore, we posit that the DL models from DeepPurpose considered herein, being sequence-based, offer better scalability. Given that only ligand and protein sequences are required for the prediction, and not all-atom 3D structures, we expect that increasingly large databases could be used for training in future iterations by running high-throughput experimental binding affinity assays (without the need for concurrent X-ray or NMR experiments) as a filtering strategy in early drug discovery. Considering that the observed performances of these models for both benchmarks of CASF-2016 and the small-scale virtual screening are better or on par with more sophisticated structure-based DL models, there is an argument to be made in favor of more scalable approaches. Additionally, we note that the public and private experimental binding affinity databases are highly fragmented, in that they are produced separately, with potentially different types of assays, or at different scales. While efforts such as the PDBbind database manage to carefully curate a high-quality single dataset, not all data sources can be incorporated because of the type of data (e.g. experimental data without associated structures being available) or because of data access (e.g. behind a private paywall). Our meta-modeling may address some of these issues by pooling only prediction outputs from models trained on diverse public or private databases with or without available structures, protecting the databases themselves if necessary and maintaining performance quality. Subsequent in-depth analysis could also identify prediction synergy and complementarity of all available tools as we demonstrated in Fig. 6. Finally, there might be value in expanding families of meta-models with different input data, each family being optimized to target a particular protein or class of proteins and/or ligands as demonstrated in Table 4.
@@
We acknowledge some limitations of our study. First, we primarily focused on a single dataset, PDBbind RefinedSet, for both training and testing. In fact, in light of recent positive results from training models on the full PDBbind General Set of ~19,000 complexes 42, 47, there is a scope for further scale-up of the model training process with larger data. It is worth mentioning, however, that when Jones et al. used the PDBbind general or refined set for model training, they observed the tradeoffs between data quality and quantity in performance of different models: SG-CNN performed better with the general set (in favor of data quantity); 3D-CNN performed better with the refined set (in favor of data quality). On the other hand, we started with DL models pre-trained on BindingDB of ~66,000 complexes and either employed them “out of the box” or fine-tuned them on the PDBbind RefinedSet (excluding the CoreSet). The two BindingDB-based meta-models, ED1 and ED2, perform better than any of the stand-alone docking or DL tools (Pearson correlation 0.62-0.64 vs. 0.47-0.59). We also tested the two PDBbind fine-tuned models, D1F and D2F, on the primary training set of BindingDB, to ascertain the degree to which the models retained their “out of the box” performance on the primary training set. We found reasonably good correlations for D1F and D2F (r = 0.636 to 0.711) (Supplementary Fig. S4) compared to the pre-trained D1 (r = 0.8) 57. However, the PDBbind-trained D3 showed poor prediction on BindingDB (r < 0.2). More importantly, while the CASF-2016 benchmark set has been well accepted, more benchmark sets of high quality would be needed to test and compare models in a more objective manner and also to challenge model generalization. Our virtual screening benchmark is also very limited and our models tend to perform well for ligands with low molecular weight. In this sense, the good performances of our meta-models do not imply model generalizability. We also note that we neither fine-tuned the docking tools nor conducted hyperparameter optimization for the DL models, which might help performance as well.

### v3

@@ Introduction
In this work, we aim to improve ligand-protein binding affinity prediction by combining both classical scoring functions, which are built upon physical force fields, and sequence-based DL models into a single integrative framework of meta-models (also called ensemble, fusion, or hybrid models). Meta-models are a type of ensemble learning by combining diverse algorithms and have been successful in machine or statistical learning for improved results with more robustness and less bias than individual algorithms (known as stacked generalization or super learning) 74, 75. Meta-models, termed blending in associated publications, were also very successful in addressing the Netflix Prize challenges 76-79. In predicting ligand-protein binding affinity, ensemble models of multiple 3D CNNs have been shown to outperform single CNNs and previous models 50. On the other hand, theory-driven physical models and data-driven ML or statistical models are complementary to each other. The former offers physical interpretability and constraints, whereas the latter is better at complex pattern discovery from massive data. For this reason, hybrid approaches have been favored in recent years 72, 73, 80-83. In general, the intuition behind meta-modeling is that different solutions to a problem could be modeling non-overlapping aspects of the ground truth. These different solutions may be trained on data with their own systematic biases, or might have architectures that access some limited subspace of the possible model space. When combined, an ensemble of such solutions might serve to better represent the ground truth. We have designed a modeling framework that allows the inclusion of structure-based classical scoring functions and sequence-based DL models, as well as explicit inclusion of covariates such as physicochemical features of ligand, proteins, or ligand-protein complexes (Figure 1 and Supplementary Figure S1). In this study, we first focus on 2 empirical scoring functions (ESFs), SMINA and Vinardo (implemented in smina) 20 together with consensus docking strategies based on experimental or predicted ligand poses. For DL models, we use sequence-based DL models of 10 different architectures with multiple training strategies, employing an open-source DL library for drug-target interaction prediction called DeepPurpose 57 and two ligand-protein binding databases, BindingDB 84 and PDBbind. We investigate both pre-trained models from the DeepPurpose library and our own fine-tuned or de novo-trained models using PDBbind, which is the main data used by the ESF models. In addition, by merging all prediction scores from repeated cross-validations for all DL models under study, we investigate top principal components from principal component analysis (PCA) as additional prediction scores. Finally, by taking all prediction scores from both ESF and DL models as well as molecular weights of ligands, we build meta-models using 4 machine learning algorithms and 11 feature sets as our final tools (44 in total) for ligand-protein binding affinity prediction.
@@ Methods
We calculate in two ways, depending on the choice of docking poses: (1) Experimental pose-RMSD filtering, where the RMSD is calculated for a ligand pose predicted by a docking method relative to the experimental structure of the ligand (when available); and (2) Consensus pose-RMSD filtering, where the RMSD is calculated between the docking poses for the same ligand-target pair as predicted by two different docking scoring methods, Vinardo and SMINA. Details of these filtering approaches are provided in Supporting Information. We also chose a cutoff of 3 Angstroms as representing a reasonably good quality match between poses based on our manual inspection of RMSD distributions (Supplementary Figure S2A). This is a more lenient cutoff relative to other analyses 20, 23 that suggest a 2-Angstrom cutoff. Our approach is a compromise between the need for higher quality matches, and the need to have sufficient numbers of data points to train our models as stricter cutoffs lead to smaller dataset sizes.
@@
Virtual Screening Benchmarks
@@
For a more general applicability, we also performed a virtual screening using an external benchmark dataset, LIT-PCBA, of active and inactive ligands for 15 target proteins 91. Although our models were neither explicitly designed nor trained to distinguish active and inactive ligands from each other (i.e., binary classification for functional activity), our assumption was that active ligands tend to have higher binding affinities (i.e., lower Kd/Ki) for target proteins than inactive ligands. This assumption is weak because there is an unclear causal relationship between binding affinity (Kd, Ki) and functional activity (IC50, EC50). We also note that we used ligand-protein complexes with Kd or Ki from both BindingDB and PDBbind for consistency. LIT-PCBA is a recently developed virtual screening dataset which was reported to be less biased than previous datasets. It is based on PubChem BioAssays which includes 4 potency types of IC50, EC50, Kd, and Ki. We point out that LIT-PCBA did not distinguish the 4 potency types from each other. As a proof-of-concept, with the aforementioned caveat in mind, we randomly selected 10 active and 10 inactive ligands for each of the 15 proteins in LIT-PCBA (i.e., a total of 300 ligand-protein pairs) to predict their binding affinities by our models. The active and inactive ligands were chosen from the combined training and validation sets in LIT-PCBA. We generated three-dimensional SDF files for each of the ligands from the SMILES strings. We also randomly selected only one of the PDB structures (in mol2 format) for each target protein in this virtual screening exercise (further details in Supporting Information). The evaluation was done in two ways: (1) comparisons of predicted binding affinity distributions between active and inactive ligands for each protein with Wilcoxon and Mann-Whitney U tests for significant differences; (2) an enrichment or fraction of active ligands in the top 5 highest-affinity ligands predicted for each protein (i.e., in the top 25% of the 20 evaluated ligands for each protein).
@@ Results
As our first group of DL models, we investigated either pre-trained (i.e., published) or de novo-trained models using the BindingDB (BDB) or PDBbind-RefinedSet (PDBb) data (model class D1, D2, or D3, respectively) (Table 2). Based on mean-squared-error (MSE), we selected the following top 6 out of the 12 de novo-trained models for subsequent analyses, which show comparable performance to the DeepPurpose pre-trained models with MSE < 0.6: CNN-CNN, Daylight-AAC, Morgan-AAC, Morgan-CNN, MPNN-CNN, and Transformer-CNN (Supplementary Figure S3). As the overlap between the published and de novo-trained models is only two, CNN-CNN and MPNN-CNN, this expanded the family of BDB-trained base models to a total of 10 with competitive performances in this study (Table 2). In terms of mean cross-validation (from 10x 5-fold CV), the best performing BDB-pre-trained, BDB-de-novo-trained, and PDBb-trained models were Daylight-AAC, Daylight-CNN, and CNN-AAC (Table 2). The model average Pearson correlation coefficients were 0.468, 0.475, and 0.746 for the CoreSet, respectively (Figure 3A; D1, D2, and D3). The two BDB-trained models perform worse than the docking tools (Figure 2). Full performance results are provided in Supplementary Tables S2 and S3. We note that the PDBb-trained models performed poorly for BDB-Kd predictions with no correlations (Supplementary Figure S4).
@@
As our second group of models, we investigated fine-tuned models using PDBb for both BDB-pre-trained and BDB-de-novo-trained models (D1F and D2F, respectively). In terms of mean cross-validation (from 10x 5-fold CV), the best performing PDBb-finetuned models for BDB-pre-trained and BDB-de-novo-trained models were CNN-CNN and MPNN-AAC. The model average Pearson correlation coefficients were 0.716 and 0.749 for the CoreSet, respectively (Figure 3A). Full performance results are provided in Supplementary Table S3. As for BDB-Kd prediction, all PDBb-finetuned models maintained good performances with Pearson correlations of 0.6–0.7, compared to predictions by the BDB-trained models (Supplementary Figures S4 and S5).
@@
As our third or final group of DL models, we investigated PCA projections of all predictions from those fine-tuned models for the BDB-pre-trained (D1F) or BDB-de-novo-trained (D2F) in the second model group (6*5*10=300 D1F or D2F model instances from 10x 5-fold CV for each model architecture) or those PDBb-trained models (D3) in the first model group (10*5*10=500 model instances). The first principal component (PC1) shows the best performance in all PCA, Pearson correlation coefficients for D1F, D2F, and D3 being 0.715, 0.749, and 0.747, respectively (Figure 3A; D1FP, D2FP, and D3P). We also performed PCA on the merged set of all predictions from D1F, D2F, and D3 (300+300+500=1100 model instances), the Pearson correlation coefficient for PC1 being 0.744 (Figure 3A; DAP). As expected, the PC1 is highly correlated with the mean prediction with the largest Pearson correlation coefficient of > 0.99 in all cases, while the second most correlated PC was not the second PC (PC2). By looking at performances of all cross-validated fine-tuned or PCA models, we observe that PC1 and D3 models tend to perform better than D1 or D2 models for the CoreSet in terms of Pearson correlation coefficients or mean squared errors (Figure 3B).
@@
Given that we have investigated the 38 DL models (6 D1, 6 D2, 10 D3, 6 D1F, 6 D2F, 1 D1FP, 1 D2FP, 1 D3P, and 1 DAP) as well as the 4 docking models, we first compared their predictions by Pearson correlation values among themselves. Figure 4A shows a heatmap of Pearson correlation values for all pairs of predictions for the CoreSet. As expected, high correlations (i.e., similar predictions) are observed among (1) the 4 docking models and (2) the fine-tuned or PC1 models (D1F, D2F, D1FP, D2FP, D3P). The BindingDB-based models (D1 and D2) are less correlated among themselves. Correlations between the experimental binding affinity and those predictions show that the top model is D1F|Daylight_AAC (pre-trained and fine-tuned) and the next top 3 models are PC1 models (Figure 4B). We also performed another comparison analysis by prediction errors or deviations from the experimental binding affinity (Supplementary Table S3). Figures 4C and 4D show a heatmap and a box-whisker plot of prediction deviations or absolute errors for the 42 models. The fine-tuned or PC1 models tend to show lower errors than the BindingDB models, which in turn tend to show lower errors than the docking models. The top 5 models in terms of median errors are the 4 PC1 models and D2F|MPNN_Transformer (de novo-trained and fine-tuned). Similar but more pronounced correlations and deviations are observed for the training set, as expected (Supplementary Figure S6).
@@
To further parse the differences between the outputs of the prediction tools, we assessed whether certain properties of the proteins or ligands are over- or under-represented in the complexes which were: (1) Best modeled by empirical docking; (2) Best modeled by deep learning; (3) Best modeled by meta-modeling. We also explored the top 50 complexes as predicted by the XGBoost Meta-model for ED-A-P to see if the performance of the meta-models relative to the experimental data was contingent on protein or ligand properties. We extracted protein features from the UniProt database and ligand features from the SwissADME database. We tested the significance of the occurrence of certain annotations (using Fisher’s exact test) and of the distributions of certain quantitative features (using a two-sided Wilcoxon test) for each of the four subsets of complexes. Specifically, we considered whether several UniProt features were annotated more or less frequently, and whether certain Gene Ontology (GO) 93, domain and pathway annotations occurred with different probabilities in the subsets. We also compared the distributions of the ligand SwissADME features. The results (Supplementary Table S9) indicate that there are practically no significant (nominal p-value ) protein features prioritized in the four cases considered. This may be due to the fact that several proteins (and their associated UniProt IDs) are duplicated in the CoreSet. The number of reviewed proteins with UniProtIDs drops from 243 to 60 after removal of all duplicates. The highly significant results with nominal p-values (both Wilcoxon and t-tests) are the distributions of several ligand features including MW, lipophilicity, and solubility measures in the structures for which the empirical docking tools perform the best. There is also some hint of significance for MW, the “#Rotatable bonds” and “#H-bond acceptors” in the structures for which the meta-models perform the best.
@@
Although the above results suggest that top-ranking high-affinity ligands predicted by our models could be prioritized as hits in virtual screening, we also tested our models using the virtual screening benchmark dataset of LIT-PCBA to distinguish 10 active ligands from 10 inactive ligands for 15 proteins (see Methods; Supplementary Table S10). As our models were not trained for binary classification of active and inactive ligands, our hypothesis was that active ligands tend to have higher affinity scores than inactive ligands. To test the hypothesis, we evaluated the fraction of active ligands in the 5 highest-scoring ligands for each protein (Table 5) and the distribution difference of predicted affinity scores between 10 active and 10 inactive ligands for each protein (Wilcoxon and Mann-Whitney U tests; Supplementary Table S11). The results show that our 5 highest-scoring ligands by either averaged DL models or averaged meta-models tend to be active ligands (i.e., >50%) for 11 out of the 15 proteins and that active ligands tend to have higher affinity scores than inactive ligands for most of the proteins. We also observe that our DL models and meta-models tend to predict higher fractions of active ligands (>=80%) than the baseline docking tools, SMINA or Vinardo (Table 5). Specifically, all the top 5 predicted ligands by D3|Transformer-CNN are active ligands for all the 15 proteins. More than 50% of the top 5 predicted ligands by most of the EW meta-models are active ligands for 13 out of the 15 proteins. We note, however, that, as in the section on prediction synergy, there is variability among the target proteins in terms of the best-performing model. Different protein targets may benefit from the use of different modeling techniques, conditional on their physicochemical properties. This is also observed in terms of the target-specific behavior of the p-values (Supplementary Table S11).
Overall, we conclude that the family of our models are competitive to several state-of-the-art structure-based models and also useful for virtual screening applications.
@@ Discussion
Overall, our framework corroborates existing work on ensemble learning, stacked generalization, super learner, and consensus-based improvements in predictions. Our machine-learning-based meta-models allow for the flexibility of adding or removing any component tools, physicochemical properties, or molecular descriptors as part of feature engineering. Models trained on multiple datasets (e.g. BindingDB and PDBbind) and on multiple data types (e.g. sequence and/or structure) may be unified to reduce individual biases and improve predictions. Furthermore, we posit that the DL models from DeepPurpose considered herein, being sequence-based, offer better scalability. Given that only ligand and protein sequences are required for the prediction, and not all-atom 3D structures, we expect that increasingly large databases could be used for training in future iterations by running high-throughput experimental binding affinity assays (without the need for concurrent X-ray or NMR experiments) as a filtering strategy in early drug discovery. Considering that the observed performances of these models for both benchmarks of CASF-2016 and the small-scale virtual screening are better or on par with more sophisticated structure-based DL models, there is an argument to be made in favor of more scalable approaches. Additionally, we note that the public and private experimental binding affinity databases are highly fragmented, in that they are produced separately, with potentially different types of assays, or at different scales. While efforts such as the PDBbind database manage to carefully curate a high-quality single dataset, not all data sources can be incorporated because of the type of data (e.g. experimental data without associated structures being available) or because of data access (e.g. behind a private paywall). Our meta-modeling may address some of these issues by pooling only prediction outputs from models trained on diverse public or private databases with or without available structures, protecting the databases themselves if necessary and maintaining performance quality. Subsequent in-depth analysis could also identify prediction synergy and complementarity of all available tools as we demonstrated in Fig. 6. Finally, there might be value in expanding families of meta-models with different input data, each family being optimized to target a particular protein or class of proteins and/or ligands as demonstrated in Tables 4 and 5.
@@
We acknowledge some limitations of our study. First, we primarily focused on a single dataset, PDBbind RefinedSet, for both training and testing. In fact, in light of recent positive results from training models on the full PDBbind General Set of ~19,000 complexes 42, 47, there is a scope for further scale-up of the model training process with larger data. It is worth mentioning, however, that when Jones et al. used the PDBbind general or refined set for model training, they observed the tradeoffs between data quality and quantity in performance of different models: SG-CNN performed better with the general set (in favor of data quantity); 3D-CNN performed better with the refined set (in favor of data quality). On the other hand, we started with DL models pre-trained on BindingDB of ~66,000 complexes and either employed them “out of the box” or fine-tuned them on the PDBbind RefinedSet (excluding the CoreSet). The two BindingDB-based meta-models, ED1 and ED2, perform better than any of the stand-alone docking or DL tools (Pearson correlation 0.62-0.64 vs. 0.47-0.59). We also tested the two PDBbind fine-tuned models, D1F and D2F, on the primary training set of BindingDB, to ascertain the degree to which the models retained their “out of the box” performance on the primary training set. We found reasonably good correlations for D1F and D2F (r = 0.636 to 0.711) (Supplementary Fig. S4) compared to the pre-trained D1 (r = 0.8) 57. However, the PDBbind-trained D3 showed poor prediction on BindingDB (r < 0.2). More importantly, while the CASF-2016 benchmark set has been well accepted, more benchmark sets of high quality would be needed to test and compare models in a more objective manner and also to challenge model generalization. Our virtual screening benchmarks are also very limited and our models tend to perform well for ligands with low molecular weight. In particular, the binary classification problem for screening of active or functional ligands would require a new model development for higher accuracy rather than repurposing exclusively affinity-based regression models as in this study. In this sense, the good performances of our fine-tuned DL models or meta-models do not imply model generalizability. We also note that we neither fine-tuned the docking tools nor conducted hyperparameter optimization for the DL models, which might help performance as well.
@@
Supporting Information
Supporting Information contains supplementary methods, supplementary figures, and supplementary references.